Shift report 20011115 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 15 2001 - 07:34:33 EST

  • Next message: Pawel Staszel: "Re: DC and Run Plan (Re: Run plan)"

    ------------------------------------------------------------
    Date: Thu Nov 15 07:34:33 2001, Shift: 00:00-08:00
    Supervisor: BHS, with: 
    Log book#10 pages 141 - 144
    
    Shift Summary:
    ------------------------------------------------------------
    1: A store had just been dumped when I came on. They got
    beam again 01.50, lasting until 05.50, and then they
    went into beam studies. This will last until 18.00 today.
    
    2: Trine completed the MRS goal at 40deg last shift, so I
    moved it to 90 deg. D5 was set to 1A field - 164 amps.
    Pedestal runs:
    5808 TPC
    5809 DVM
    5810 Globals
    
    Run 5811 is a <10min data run with some 40k trig 6 - I found
    that T3s field 2 had tripped or not come up properly. 
    Badtrip did not notice this, though... 
    
    Run   Trig 4   Trig 6     Setting
    5812    191k     104k     MRS@90, D5@1A (164 amps)
                              FS@12, D1-D4@0.5A
    5813    200k     107k     Same
    5814    345k     106k     D5@1B, otherwise the same 
    
    3: No problems except for the one T3 trip.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 15 2001 - 07:35:10 EST